### v1

# Strategies for the alignment of electronic states in quantum-dot tunnel-injection lasers and their influence on the emission dynamics

Abstract
In quantum-dot tunnel-injection lasers, the excited charge carriers are efficiently captured from the bulk states via an injector quantum well and then transferred into the quantum dots via a tunnel barrier. The alignment of the electronic levels is crucial for the high efficiency of these processes and especially for the fast modulation dynamics of these lasers. In particular, the quantum mechanical nature of the tunneling process must be taken into account in the transition from two-dimensional quantum well states to zero-dimensional quantum dot states. This results in hybrid states, from which the scattering into the quantum-dot ground states takes place. We combine electronic state calculations of the tunnel-injection structures with many-body calculations of the scattering processes and insert this into a complete laser simulator. This allows us to study the influence of the level alignment and limitations due to inhomogeneous quantum-dot distributions. We find that the optimal alignment deviates from a simple picture in which the of the quantum-dot ground state energies are one LO-phonon energy below the injector quantum well ground state.
I Introduction
Tunnel-injection (TI) lasers with quantum dots (QD) as active material promise improved emission properties, in particular faster modulation speed in optical data transmission networks. Sun et al. (1993); Bhattacharya and Ghosh (2002); Han and Asryan (2008); Bhowmick et al. (2014); Gready and Eisenstein (2010) In conventional QD lasers with current injection pumping, the pump process generates excited carriers in the bulk barrier states. Capturing these carriers into the QD states requires substantial energy dissipation via carrier-phonon-scattering in combination with carrier-carrier scattering. This limits the generation efficiency of excited carriers in the QD states and slows down the modulation speed of these lasers. The TI-laser design utilizes an injector quantum well (QW) that is directly coupled to the QDs via a tunnel barrier. The injector QW provides a much larger scattering cross section to the excited carriers driven by the carrier injection, thereby enhancing the capture rate of carrier from the bulk states. Tunnel coupling facilitates efficient transfer of excited carriers from the injector QW into the QD states.
The design of TI-QD lasers requires proper alignment of the electronic states between the injector QW and the QDs to exploit the advantages of this system. The dominant carrier scattering process, which contributes to the tunnel-assisted carrier transfer from the injector QW to the QDs has been identified as carrier interaction with LO phonons. Sun et al. (1993); Bhattacharya and Ghosh (2002); Bhowmick et al. (2014); Gready and Eisenstein (2010); Michael et al. (2018); Lorke et al. (2018, 2022) Hence the original design goal was to engineer the structure in a way the QD ground states are positioned one LO-phonon energy below the band edge of the injector QW.
A closer look reveals two challenges. Due to an inhomogeneous distribution of QD confinement energies, only some of the QDs have optimal tunnel coupling. Moreover, the tunnel coupling does not occur from the bottom of the injector QW. In Refs. Michael et al. (2018); Lorke et al. (2018, 2022) it was shown that the quantum mechanical coupling of the 2d and 0d subsystems has to be taken into account. It leads to hybrid states which are partly localized in the in the injector quantum well as well as in the QDs. The more elaborate picture for the transfer of excited carriers from the QW into the QD thus involves carrier scattering processes from such a hybrid state into the QD ground state through emission of LO phonons.
We have developed microscopic models for the description of the excited carrier transfer from the injector QW to an inhomogeneously distributed QD ensemble. Starting from electronic state calculations for the coupled 2d-0d system, interaction matrix elements for the carrier-phonon interaction are determined. These are included in a quantum kinetic calculation of scattering rates, which account for polaron energy renormalizations and deviations from a strict energy conservation in terms of free-particle states for the scattering processes. Seebeck et al. (2005); Michael et al. (2018); Lorke et al. (2018, 2022) The results determine capture rates of excited electrons into the QDs as function of their detuning and the resulting carrier dynamics.
In this paper we integrate our model for the carrier kinetics in tunnel-coupled injector QW and QD system into a complete laser simulator. The latter includes a spatially resolved description of the carrier transport in a structure with 6 injector QWs and the coupling of the excited carrier recombination in the QDs via spontaneous and stimulated emission to a self-consistent laser model.
II Theory
We consider a QD ensemble with size and composition fluctuations leading to a Gaussian distribution of ground state (s-shell) energies with a FWHM of 15meV. The maximum of the distribution is centered at an interband transition energy of 800meV (1550nm). The situation is sketched in Fig. 1 with the blue line. Fig. 1 shows three examples of the
Figure 1: Inhomogeneous distribution of QD ground state energies (blue line) and hybridized p-shell states within the injector QW continuum (shaded area). Top: Examples for three different QDs with zero-dimensional ground state (s-shell). These QDs do not have a localized p-shell, as these states hybridize with the injector QW continuum states.
QDs in the inhomogenoues ensemble with their ground state energy (solid colored line) and the associated hybridized injector QW states (colored shaded areas). These are no longer zero-dimensional states (associated with the former p-shell) Lorke et al. (2022), but have a two-dimensional component in the injector QW due to the tunnel coupling. Due to the differences in energetic alignment between QDs in the inhomogeneous ensemble, they couple differently to the injector QW.
In this work, we compare the current TI design, where the energetic alignment between the QD ground state and the injector QW ground state is tuned to the LO phonon energy (Alignment A in what follows), with a situation that we propose in this work, where the injector QW is shifted to lower energies in an effort to optimize the scattering efficiency (Alignment B).
To study switch-on and modulation properties of lasers based on TI structures as active material, we set up a model that includes both the microscopic carrier dynamics in the TI structure on the one hand, as well as the carrier transport though the bulk material on the other.
The carrier transport in the bulk material is described by
$\frac{d}{dt}N_{\text{res}}(z,t)|_{\text{transport}}=\frac{d}{dz}\left[D_{N}\frac{d}{dz}N_{\text{res}}(z,t)+\mu_{N}N_{\text{res}}E(z,t)\right]$ (1)
with boundary condition
$J_{N}=q\left[D_{N}\frac{d}{dz}N_{\text{res}}(0,t)+\mu_{N}N_{\text{res}}E(0,t)\right]\leavevmode\nobreak\ .$ (2)
In this equation $N_{res}$ is the three-dimensional carrier density of the bulk material, $D_{N}$ and $\mu_{N}$ are the diffusion coefficient and mobility of the carriers and $E(z,t)$ is the internal field obtained by a solution of Poisson’s equation for the carrier density.
The carriers in the bulk material are couples to the active region, consisting of 6 QW-QD layers that couple to the bulk carrier density
$\begin{split}\frac{d}{dt}N_{\text{res}}(z,t)&=\frac{d}{dt}N_{\text{res}}(z,t)|_{\text{transport}}-\frac{N_{\text{res}}(z,t)}{\tau_{res}}\\ &-N_{\text{res}}(z,t)\sum_{i=1}^{M}\frac{N_{QD}^{i}}{\tau^{i}_{cap}}(1-f^{i;e,h})\\ &+\sum_{i=1}^{M}\frac{N_{QD}^{i}}{\tau^{i}_{\text{esc}}}f^{i;e,h}\left(1-\frac{N_{\text{res}}}{D_{res}}\right)\\ &-\sum_{j}\frac{N_{\text{res}}}{\tau_{cap,IW}}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\\ &+\sum_{j}\frac{N^{j}_{IW}}{\tau_{esc,IW}}\left(1-\frac{N_{\text{res}}}{D_{res}}\right)\leavevmode\nobreak\ .\end{split}$ (3)
Here, $D_{res}$ ($D_{IW}$) is the density of states of the bulk (injector QW), $N^{j}_{IW}$ the injector QW carrier density of layer $j$, $f^{i;e,h}$ the electron and hole occupation of the of the QD s-shell in the i-th sub-ensemble. $\tau^{i}_{cap,esc}$ is the capture (escape) scattering time from the bulk into the QD and $\tau_{cap,esc,IW}$ is the capture (escape) scattering time between bulk and injector QW.
The equation of motion for the injector QW density follows as
$\begin{split}\frac{d}{dt}N^{j}_{IW}&=\frac{N^{j}_{IW}}{\tau_{IW}}-\frac{N^{j}_{IW}}{\tau_{\text{tun}}}\sum_{i=1}^{M}N_{QD}^{i}(1-f^{i;e,h})\\ &+\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\sum_{i=1}^{M}2N_{QD}^{i}\frac{f^{i;e,h}}{\tau_{tun,b}}\\ &+\frac{N_{\text{res}}(z_{W},t)}{\tau_{cap.IW}}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\\ &-\frac{N^{j}_{IW}}{\tau_{esc,IW}}\left(1-\frac{N_{\text{res}}(z_{W},t)}{D_{res}}\right)\leavevmode\nobreak\ ,\end{split}$ (4)
with $\tau_{tun}$ is the tunnel scattering time between the hybridized injector QW state and the QD s-shell, as describe in our previous publication Lorke et al. (2022).
The QD populations $f^{i;e,h}$ follow from the semiconductor Bloch equations Haug and Koch (2004)
$\begin{split}\frac{d}{dt}f^{i;e,h}=&-\gamma_{c}f^{i;e,h}+\frac{N_{\text{res}}(z_{D},t)}{2\tau_{c}ap}(1-f^{i;e,h})\\ &-\frac{f^{i;e,h}}{\tau^{i}_{\text{esc}}}(1-\frac{N_{\text{res}}(z_{D},t)}{D_{res}})\\ &-2\text{Im}\left(\Omega_{\text{lase}}\psi^{i,\ast}\right)+\sum\limits_{j}\frac{N^{j}_{IW}}{2\tau_{\text{tun}}}(1-f^{i;e,h})\\ &-\sum\limits_{j}\left(1-\frac{N^{j}_{IW}}{D_{IW}}\right)\frac{f^{i;e,h}}{\tau_{tun,b}}\end{split}$ (5)
with the interband polarization
$\frac{d}{dt}\psi^{i}=-(i\omega+\gamma_{h})\psi^{i}-i\frac{\Omega_{\text{lase}}}{\hbar}\left(1-f^{i;e}-f^{i;h}\right)\leavevmode\nobreak\ ,$ (6)
where $\gamma_{h}$ is the homogeneous linewidth.
III Results
We consider an inhomogeneous QD ensemble with size and composition fluctuations, leading to a Gaussian distribution of ground state (s-shell) energies. The QD layers are separated from the injector QW by a thin 2.1 nm thick tunnel barrier. The active material of the laser consists of 6 QW-QD layers separated by 10.5 nm wide spacers layers.
Figure 2: Inverse tunnel scattering time as a fucntion of QD s-shell energy for the present experimental alignment (red line) to the QD distribution (black line) as well as the optimized alignment (blue line).
Previously Lorke et al. (2022), we have demonstrated that current TI designs (alignment A) couple rather inefficiently to the injector QW. This is shown (red line) in Fig. 2 together with the QD distribution (black line). Here, the dashed line refers to the situation, where the p-shell is hybridized, while for the dotted line the p-shell is confined and the d-shell hybridizes with the QW. For details, see Ref. Lorke et al. (2022). In comparison our proposed alignment B (blue line in Fig. 2) has an improved carrier scattering at the maximum of the QD distribution. Such an alignment could be generated by, e.g., modifying the injector QW width or concentration profile. One of the main reasons for the improved scattering rates in alignment B is that the most efficient carrier scattering does not occur from the bottom of the QW but from higher lying states due to symmetry considerations. Hence, even in a independent particle picture the QW needs to be tuned accordingly.
In Fig. 3 input-output characteristics are shown for alignment A and B. Notably, the proposed alignment B has a higher threshold current than the current alignment A. The reason is that carriers are more evenly distributed over the 6 active layers with the optimized alignment (see Fig. 4), leading to a higher carrier density to be accumulated in the device before lasing occurs. This is caused by the additional carriers in the injector QWs of the outer layers, that need to be populated for carrier scattering into the QD ground states to occur.
Figure 3: Input/Output characteristics for TI laser devices with 6 active layers and current (red) and optimized (blue) alignment between the injector QW and the QD ensemble
To investigate this further, the total carrier density $(N_{res}+N_{IW}+N_{QD}f^{i;e})$ is shown in Fig. 4 for a pump power 10% above the laser threshold for both devices. Electrons (Holes) are generated by the contact in this Figure at the left (right) edge of the device. For the device with the current experimental alignment A, the slower tunnel scattering rate leads carriers more localized in the central layers of the device.
Figure 4: total electron density as a function of position and time for the current (blue) and optimized (red) alignment situation.
While the higher threshold seems to indicate that the proposed "optimized" alignment B performs worse, the situation is reversed when investigating dynamical laser properties such as the switch-on process. The dynamic of the output power is shown in Fig. 5. For better visibility, only the first 300 ps are shown, whereas the simulation is run to reach steady state.
Figure 5: Output power as a function of time for a pump power 10% above the laser threshold for the current alignment (red) and the optimized alignment (blue)
As the carrier spread out more evenly over the 6 layer structure, the onset of lasing shows a small delay of about 10ps Clearly, however, the optimized structure shows a significantly faster switch-on process. Comparing the results to laser rate equation models Coldren and Corzine (1995) reveals that the optimized structure should be close to critical damping of the relaxation oscillations Lorke et al. (2011), as only half an oscillation cycle is visible. This finding already leads us to expect enhanced modulation capabilities of the optimized structures in comparison with current experimentally realized ones.
IV Acknowledgment
This work was partially supported by The Israel Science Foundation under grant number 460/21. I.K. acknowledges the support of the Rothschild Fellowship (The Yad Hanadiv Foundation), and of the Helen Diller Quantum Center at Technion. M.L and F.J acknowledge support from the German Science Foundation (DFG) under grant number JA619/16-2 and CPU time at HLRN (Berlin/Göttingen).